OpenStudy (anonymous):

Which are the solutions to the equation |2x + 3| - 5 = 12?

OpenStudy (yb1996):

First you keep isolate the absolute value |2x + 3| = 17 Then you make two equations and solve. The first equation is: 2x + 3 = 17 2x = 14 x = 7 The second equation would be: 2x + 3 = -17 2x = -20 x = -10

OpenStudy (yb1996):

So the solutions are x=7 and x= -10